### v2

@@ Abstract
This paper introduces innovative enhancements to the K-means and K-nearest neighbors (KNN) algorithms based on the concept of Gini prametric spaces. Unlike traditional distance metrics, Gini prametrics incorporate both value-based and rank-based measures, offering robustness to noise and outliers. The paper’s main contributions include (1) proposing a Gini prametric that captures rank information alongside value distances, (2) presenting a Gini K-means algorithm that is proven to converge and demonstrates resilience to noisy data, and (3) introducing a Gini KNN method that rivals state-of-the-art approaches like Hassanat’s distance in noisy environments. Experimental evaluations on 14 datasets from the UCI repository reveal the superior performance and efficiency of Gini-based algorithms in clustering and classification tasks. This work opens new avenues for leveraging rank-based prametrics in machine learning and statistical analysis.
@@ 1 Introduction
Machine learning, data analysis, and pattern recognition are mainly divided into two fields corresponding to two different tasks. The former is employed for supervised classification (with labeled data), while the latter is used for unsupervised classification or clustering (with unlabeled data). In each field, numerous algorithms are available based on different metrics, however some distance metrics are used in both contexts, classification and clustering. For instance, K-nearest neighbors or KNN Fix & Hodges (1951); Cover & Hart (1967) and K-means algorithms Steinhaus (1957); Lloyd (1957; 1982); Ball & Hall (1965); MacQueen (1967) were, at the moment of their creation, based on the Euclidean distance, before moving to a more flexible metric i.e. the Minkowski -distance (Oti et al., 2021).
@@
On the one hand, K-means involves assigning points to clusters by iteratively measuring the distance between the points and the centroids of the clusters, with centroids being updated at each step of the algorithm until convergence. Jain (2008; 2010) shows that more than one thousand algorithms are available for K-means clustering. The developments of K-means are either concerned with initialization, convergence and partitioning the data (see for instance Pelleg & Moore (1999); Steinbach et al. (2000)) or the distance metric to be used to measure the proximity or the similarity between points and centroids, see for example the Kernel K-means to incorporate non-linearity in distances Schölkopf et al. (1998), the Fuzzy C-means Eschrich et al. (2003), the K-menoids based on the median Kaufman & Rousseeuw (2005), the X-means with Akaike or Bayesian information criteria Pelleg & Moore (2000), and recently Fair K-means Bateni et al. (2024) including fairness in distance functions.
@@
On the other hand, KNN allows assigning points of the testing dataset to labels by measuring the distance between these test points and the K labeled points in the training dataset that are closest. The algorithm is therefore dependent on the number of neighbors K, but also, and more importantly, dependent on the choice of the distance Syriopoulos et al. (2023). Alfeilat et al. (2019) show that non-convex distances are well suited for managing noise in data whereas Vincent & Bengio (2001) show that convex distances defined on non-linear manifolds may be a good strategy to refine the classification. As in the K-means, weighting the points differently Gini (1912) or by means of fuzzification may improve the accuracy of the algorithm Derrac et al. (2014). However, as mentioned in Alfeilat et al. (2019), comparing 44 distances brings out 8 distances that outperform the others on many datasets (see Section 2 for details).
@@
Prametric (or premetric) spaces have been introduced by Arkhangelskii & Pontryagin (1990). From our knowledge, apart from the use of non-convex distances, prametrics have not been employed in classification. A prametric space is a topological space that is more general than metric spaces, requiring neither symmetry nor indistinguishability, nor the validity of triangular inequality. It is provided with a function that is more general than a “distance” in the usual sense, since must satisfy non-negativity and . is then said to be a “prametric”, and since prametrics are not metrics (they usually do not satisfy the triangle inequality), they do not induced metric spaces as usual in topology. Indeed, the topology of a space induced by a prametric is necessarily sequential.
@@
The Gini index is often used in multivariate statistics because it is a rank-dependant statistics robust to noise and outliers. This robustness holds true for many tasks such as regressions Yitzhaki & Schechtman (2004); Yitzhaki & Lambert (2013) or compression Charpentier et al. (2021). We shown in this paper that a prametric Gini can be defined with an hyper-parameter and employed for supervised classification and clustering to deal with noisy data. The main contributions of this paper are threefold:
@@
3. To show that Gini KNN may compete with the KNN based on the Hassanat distance, which is considered to be the most robust to noise.
@@ 2 State of the art
Metric distances play a central role in both supervised and unsupervised learning methods. The choice of the distance is challenging in practice, because of their variety and also because of the difficulty of determining in advance which distance is the most appropriate for a given dataset. However, Alfeilat et al. (2019) demonstrate that two distances outperform other ones (among a set of 44 distances) when dealing with noisy datasets and the K-nearest neighbor (KNN) classifier: the Hassanat distance and distances based on the L1 distance (Manhattan or city block distance), such as the Lorentzian and the Canberra distances. Other studies on K-means show that robust clustering may be performed thanks to weighted Manhattan distances An & Jiang (2024), quantum Manhattan distances Wu et al. (2022).
A distance satisfies the following properties: non-negativity, symmetry, nullity if and only if the points are equal, and finally the triangle inequality. Apart from Elkan (2003) who shows that the triangle inequality may serve to accelerate the K-means algorithm, there is no justification for all four of these properties to be satisfied at the same time. It might be possible to resort to pseudo-distances or prametrics to look for more flexible functional forms. From the review made by Alfeilat et al. (2019), eight families of distances can be identified in the literature. Let be two points, we provide in Table 1 the eight families with one member of each family (or two) that receives attention from the literature.
@@
| Squared Chi squared | |
@@
In particular, Alfeilat et al. (2019) show that Hassanat’s distance Hassanat (2014) outperforms 43 other distances of the eight families based on the average precision and recall over many datasets of the UCI machine learning repository. They perform the same experiment with noise (from 10% to 90% of noise in the data with increment of 10%) and find that the same ranking is obtained in terms of averaged recalls and precisions: 1) Hassanat, 2) Lorentzian (F2), 3) Canberra (F2). The F1 and F3 distance families are less efficient, particularly the Euclidean and Cosine distances, as well as the Pearson distance (in F8) based on Pearson’s correlation coefficient.
@@ 3.1 Variance, Covariance, Norm and Distance
Nevertheless, it is well established that those measures are sensitive to outliers. A classical strategy is then to move from the L2 norm () to the L1 norm (). E.g., instead of using the average (solution of ), one could consider the median (solution of ). And the L1 norm is related to ranks, since, if and denotes a real-valued random variable with c.d.f. , while . Then
@@
when the empirical e.c.d.f. is based on ranks (if is ordered, where ), where denote the rank of among . The use of ranks has been intensively considered in statistics to avoid the influence of outliers, as discussed in Kendall (1948); Hollander (1973); Huber (1981). Hence, Spearman (1904) suggested to use Spearman’s covariance, where ranks are considered, instead of numerical values.
3.2 Gini Index and Gini Covariance
The popular Gini index is defined as a normalized version of “Gini Mean Difference”, Gini (1912), defined as the expected absolute difference between two randomly drawn observations from the same population. The later can be rewritten with the covariance operator,
And Schechtman & Yitzhaki (1987) introduced a Gini Covariance operator, defined, for a pair of variables, as
where that is centered, and therefore
@@ 3.2 Gini Index and Gini Covariance
which is a positive matrix (see Dang et al. (2019); Charpentier et al. (2021)). Thus, in dimension 2, we can define the Gini-covariance between vectors in as,
We can now use those dissimilarity measures on variables to define a prametric distance on the set of individuals.
@@ 3.3 Gini Prametric
In practice, the average rank method is usually employed to deal with ties between many points, in this case the arithmetic mean of their ranks is computed. This avoids points with same values to be weighted by different ranks (this would imply for instance a bias in the Gini index, see Yitzhaki & Schechtman (2013)).
@@
Accordingly, a strategy for computing conditional ranks is necessary for train-test splits (this will be described in Section 4). The main properties of the Gini prametric are the following.
Proposition 3.2.
@@
Straightforward from the properties of the Gini covariance (Yitzhaki & Schechtman, 2013). ∎
Linear invariance postulates that if the elements in are increasing while their ranks remain constant, then the Gini prametric is invariant to such a translation. On the other hand, if a new point is included in , or if is transformed in a such a way that the ranks increase by the same amount, the Gini prametric remains invariant. Finally, it is non-negative and symmetric, therefore following the terminology employed by Arkhangelskii & Pontryagin (1990), it can be referred to as a Gini symmetric, such that is a Gini pramatric space.
Notice that if denotes descending ranks
@@ 3.4 The Generalized Gini Prametric
In a series of papers Schechtman & Yitzhaki (1987; 2003); Yitzhaki & Schechtman (2013) define a generalized version of the Gini Mean Difference:
with the decumulative distribution function (i.e., ). Yitzhaki & Schechtman (2013) show that the Gini covariance operator is a robust version of the standard covariance operator. The robustness to outliers allows Gini-based measures to be enough flexible thanks to the hyper-parameter . This flexibility yields robust estimators of regression parameters as well as correlation coefficients Yitzhaki & Lambert (2013). Additionally, in the case of for multivariate normal data, least squares regressions become a particular case of Gini regressions. Generalized Gini principal components analysis, as introduced by Charpentier et al. (2021), enables the influence of outliers to be minimized and helps find the best sub-space to fit the data according to the value of (for and multivariate normal data, the usual principal component analysis is recovered). Following Schechtman & Yitzhaki (2003), the use of decumulative ranks raised to the power of allows the Gini prametric to be rewritten as follows
Definition 3.3.
@@
Straightforward from the properties of the Gini covariance (Yitzhaki & Schechtman, 2013). ∎
@@
The two first terms of the previous relation correspond to the two Gini covariances (up to ): with the rank of and conversely with the ranks of . The two last terms are (up to ) generalized Gini Mean Difference, which correspond to the intrinsic variability of and respectively. Everything happens as if the generalized Gini prametric measured on points belonging to different groups captures the two Gini covariances between groups in excess of the covariances within groups. In this respect, it could be included in the family of distances (F3) based on scalar products.
@@ 4.1 Gini KNN
See Appendix A.1. ∎
@@ 4.2 Gini K-means
The Gini K-means algorithm based on the generalized Gini prametric may be implemented if two requirements are met. First, convergence must be ensured for any given dataset and hyper-parameters .
Proposition 4.2.
The Gini K-means algorithm based on the generalized Gini prametric is convergent, for all .
@@
See Appendix A.2. ∎
Second, since convergence is always guaranteed, an optimal value may be derived. Setting training and testing data as , a grid search may be conducted as follows. The Gini K-means algorithm is trained on and the silhouette criterion (Rousseeuw, 1987) is computed for each value of (not exceeding 6, otherwise too much weight is placed on minimal values of ). The optimal hyper-parameter is the value that maximizes the silhouette score. Once is determined, the labels of the testing data can be inferred.
@@
Finally, other clustering techniques relying on the Euclidean distance, close to K-means, may be adapted to the generalized Gini prametric, such as the agglomerative clustering (not detailed here). An illustration is depicted in Figure 2 on the iris dataset.111 See Github.
@@ 5 Experiments
K-means and KNN classifications have been performed adapting the scikit-learn library to the generalized Gini prametric. An additional illustration is also conducted with agglomerative clustering. These algorithms are compared on the basis of 12 distances issued from the 8 families (described in Table 1): L Minkowski distance (Euclidean, Minkowski, Manhattan), L1 distance (Canberra), Inner product distance (Cosine), Squared Chord distance (Hellinger distance), Squared L2 distance (Pearson ), Shannon entropy distance (Jensen-Shannon), Vicissitude distance (Vicis Symmetric), and other distances (Hassanat). The comparison is performed over 16 datasets of UCI repository222UCI datasets repository link., without normalization in order to observe the impact of outliers (if present) on the performance. Following the study of Alfeilat et al. (2019), the 16 datasets have been chosen with regard to the excellent performance of some distances robust to outliers, in particular the Hassanat distance. The datasets are described in Table 2. The detailed metrics of the Gini K-means and Gini KNN experiments, analyzed in the following subsections, are relegated in the Appendix.
@@ 5.1 KNN experiments
The generalized Gini prametric is implemented with the supervised KNN algorithm using the scikit-learn library. To compare the 12 distances listed above, precision and recall metrics are used on the same UCI datasets described in Table 2. To identify the optimal number of neighbors, is looped from to for each distance. A grid search is then performed to select the optimal hyper-parameter of the generalized Gini prametric that maximizes the precision. Then, recall and precision are computed and a 3-fold cross-validation is performed to validate the results (see Tables 16 and 12 in Appendix A, and in Appendix and B).
@@
With 5% and 10% Noise. The experiment involves 5% and 10% noise into the data. Specifically, Gaussian noise with a mean of 0 and a standard deviation of 1 is added. This setup is designed to evaluate the robustness of the KNN algorithms under less favorable conditions, where the generalized Gini prametric is expected to outperform the other methods. In the first case, the Gini KNN achieves the highest average precision (11 top-ranked positions out of 16) and recall (10 top-ranked positions). In the 10% contamination case, the performance are much more variable. The generalized Gini prametric achieves only 6 best positions for precision and 5 for recall. Then, it comes Hassanat, Minkowski and Cosine (2 top positions for precision), and Hassanat again for recall (3 top positions), see Tables 2 to 6 in Appendix B.

### v3

@@ Abstract
This paper introduces innovative enhancements to the K-means and K-nearest neighbors (KNN) algorithms based on the concept of Gini prametric spaces (as opposed to metric spaces with standard distance properties). Unlike traditional distance metrics, Gini prametrics incorporate both value-based and rank-based measures, offering robustness to noise and outliers. The main contributions of the paper include (1) proposing a Gini prametric that captures rank information alongside value distances, (2) presenting a Gini K-means algorithm that is proven to converge and demonstrates resilience to noisy data, and (3) introducing a Gini KNN method that rivals state-of-the-art approaches like Hassanat’s distance in noisy environments. Experimental evaluations on 16 datasets from the UCI repository reveal the superior performance and efficiency of Gini-based algorithms in clustering and classification tasks. This work opens new directions for rank-based prametrics in machine learning and statistical analysis.
123
@@ 1 Introduction
Machine learning, data analysis, and pattern recognition are closely related fields, each encompassing a variety of tasks and approaches. Among the most common tasks are supervised classifications, where models are trained using labeled data, and unsupervised learning, such as clustering, which deals with finding patterns in unlabeled data. In each field, numerous algorithms are available based on different metrics, however some distance metrics are used in both contexts, classification and clustering. For instance, K-nearest neighbors or KNN (Fix and Hodges [15], Cover and Hart [10]) and K-means algorithms (Steinhaus [38], Lloyd [25], Ball and Hall [5], MacQueen [26]) were, at the moment of their creation, based on the Euclidean distance, before moving to a more flexible metric i.e. the Minkowski -distance (Oti et al. [28]).
@@
On the one hand, K-means involves assigning points to clusters by iteratively measuring the distance between the points and the centroids of the clusters, with centroids being updated at each step of the algorithm until convergence. Jain [20, 21] shows that more than one thousand algorithms are available for K-means clustering. The developments of K-means are either concerned with initialization, convergence and partitioning the data (see for instance Pelleg and Moore [29], Steinbach et al. [37]) or the distance metric to be used to measure the proximity or the similarity between points and centroids, see for example the Kernel K-means to incorporate non-linearity in distances (Schölkopf et al. [34]), the Fuzzy C-means (Eschrich et al. [14]), the K-medoids based on the median (Kaufman and Rousseeuw [22]), the X-means with Akaike or Bayesian information criteria (Pelleg and Moore [30]), and recently Fair K-means (Bateni et al. [6]) including fairness in distance functions.
@@
On the other hand, KNN allows assigning points of the testing dataset to labels by measuring the distance between these test points and the K labeled points in the training dataset that are closest. The algorithm is therefore dependent on the number of neighbors K, but also, and more importantly, dependent on the choice of the distance (Syriopoulos et al. [39]). Alfeilat et al. [1] show that non-convex distances are well suited for managing noise in data whereas Vincent and Bengio [40] show that convex distances defined on non-linear manifolds may be a good strategy to refine the classification. As in the K-means, weighting the points differently (Gini [16]) or by means of fuzzification may improve the accuracy of the algorithm (Derrac et al. [12]). However, as mentioned in Alfeilat et al. [1], comparing 44 distances brings out 8 distances that outperform the others on many datasets (see Section 2 for details).
@@
Prametric (or premetric) spaces have been introduced by Arkhangelskii and Pontryagin [3]. From our knowledge, apart from the use of non-convex distances, prametrics have not been employed in classification. A prametric space is a topological space that is more general than metric spaces, requiring neither symmetry nor indistinguishability, nor the validity of triangular inequality. It is provided with a function that is more general than a “distance” in the usual sense, since must satisfy non-negativity and . is then said to be a “prametric", and since prametrics are not metrics (they usually do not satisfy the triangle inequality), they do not induce metric spaces as usual in topology. Indeed, the topology of a space induced by a prametric is necessarily sequential.
@@
The Gini index is often used in multivariate statistics because it is a rank-dependent statistics robust to noise and outliers. This robustness holds true for many tasks such as regressions (Yitzhaki and Schechtman [44], Yitzhaki and Lambert [43]) or compression (Charpentier et al. [8]). We show in this paper that a prametric Gini can be defined with a hyper-parameter and employed for supervised classification and clustering to deal with noisy data. The main contributions are:
@@
3. To show that Gini KNN may compete with KNN based on Hassanat distance, which is considered to be the most robust to noise.
@@ 2 State of the art
Metric distances play a central role in both supervised and unsupervised learning methods. The choice of the distance is challenging in practice, because of their variety and also because of the difficulty of determining in advance which distance is the most appropriate for a given dataset. However, Alfeilat et al. [1] demonstrate that two distances outperform other ones (among a set of 44 distances) when dealing with noisy datasets and the K-nearest neighbor (KNN) classifier: the Hassanat distance and distances based on the L1 norm (Manhattan or city block distance), such as the Lorentzian and the Canberra distances. Other studies on K-means show that robust clustering may be performed thanks to weighted Manhattan distances (An and Jiang [2]), quantum Manhattan distances (Wu et al. [41]).
@@
| Pearson Chi squared | |
| Squared Chi distance | |
@@
A distance satisfies the following properties: non-negativity, symmetry, nullity if and only if the points are equal, and finally the triangle inequality. Apart from Elkan [13] who shows that the triangle inequality may serve to accelerate the K-means algorithm, there is no justification for all four of these properties to be satisfied at the same time. It might be possible to resort to pseudo-distances or prametrics to look for more flexible functional forms. From the review made by Alfeilat et al. [1], eight families of distances can be identified in the literature. Let be two points, we provide in Table 1 the eight families with one member of each family (or two) that receives attention from the literature.
In particular, Alfeilat et al. [1] show that Hassanat’s distance [17] outperforms 43 other distances of the eight families based on the average precision and recall over many datasets of the UCI machine learning repository. They perform the same experiment with noise (from 10% to 90% of noise in the data with increment of 10%) and find that the same ranking is obtained in terms of averaged recalls and precisions: 1) Hassanat (F8), 2) Lorentzian (F2), 3) Canberra (F2). The F1 and F3 distance families are less efficient, particularly the Euclidean and Cosine distances, as well as the Pearson distance (F8) based on Pearson’s correlation coefficient.
@@ 3.1 Variance, Covariance, Norm and Distance
Nevertheless, it is well established that those measures are sensitive to outliers. A classical strategy is then to move from the L2 norm to the L1 norm . E.g., instead of using the average (solution of ), one could consider the median (solution of ). And the L1 norm is related to ranks, since, if and denotes a real-valued random variable with c.d.f. , while . Then
@@
when the empirical e.c.d.f. is based on ranks (if is ordered, where , and where denotes the rank of among . The use of ranks has been intensively considered in statistics to avoid the influence of outliers, as discussed in Kendall [23], Hollander [18], Huber [19]. Hence, Spearman [36] suggested to use Spearman’s covariance, where ranks are considered, instead of numerical values. More recently, Carcea and Serfling [7] and Shelef and Schechtman [35] show the robustness of the Gini-covariance (see subsection 3.4) that can be used for the Gini prametric.
3.2 Gini Index and Gini-Covariance
The popular Gini index is defined as a normalized version of “Gini Mean Difference”, Gini [16], defined as the expected absolute difference between two randomly drawn observations from the same population. The later can be rewritten with the covariance operator,
Schechtman and Yitzhaki [32] introduced a Gini-covariance operator, defined, for a pair of variables, as
where is centered, and therefore
@@ 3.2 Gini Index and Gini-Covariance
which is a positive matrix (see Dang et al. [11], Charpentier et al. [8]). Thus, in dimension 2, we can define the Gini-covariance between vectors in as,
We can now use those dissimilarity measures to define a prametric distance.
@@ 3.3 Gini Prametric
In practice, the average rank method is usually employed to deal with ties between many points, in this case the arithmetic mean of their ranks is computed. This avoids points with same values to be weighted by different ranks (this would imply for instance a bias in the Gini index, see Yitzhaki and Schechtman [45]).
@@
Accordingly, a strategy for computing conditional ranks is necessary when new instances are added to the dataset (for train-test splits this will be described in Section 4). The main properties of the Gini prametric are the following.
Proposition 1.
@@
Straightforward from the properties of the Gini-covariance (Yitzhaki and Schechtman [45]). ∎
Linear invariance postulates that if the elements in are increasing while their ranks remain constant, then the Gini prametric is invariant to such a translation. On the other hand, if a new point is included in , or if is transformed in such a way that the ranks increase by the same amount, the Gini prametric remains invariant. Finally, it is non-negative, symmetric and does not satisfy the triangle inequality, therefore following the terminology employed by Arkhangelskii and Pontryagin [3], it can be referred to as a Gini symmetric, such that is a Gini pramatric space.
Notice that if denotes descending ranks,
@@ 3.4 The Generalized Gini Prametric
In a series of papers Schechtman and Yitzhaki [32, 33], Yitzhaki and Schechtman [45] define a generalized version of the Gini Mean Difference:
with the decumulative distribution function of (i.e., ). Yitzhaki and Schechtman [45] show that the Gini-covariance operator is a robust version of the standard covariance operator. The robustness to outliers allows Gini-based measures to be enough flexible thanks to the hyper-parameter . This flexibility yields robust estimators of regression parameters as well as correlation coefficients [43]. Additionally, in the case of for multivariate normal data, least squares regressions become a particular case of Gini regressions. Generalized Gini principal components analysis, as introduced by Charpentier et al. [8], enables the influence of outliers to be minimized and helps find the best sub-space to fit the data according to the value of (for and multivariate normal data, the usual principal component analysis is recovered). Following Schechtman and Yitzhaki [33], the use of decumulative ranks raised to the power of allows the Gini prametric to be rewritten as follows:
Definition 2.
@@
Straightforward from the properties of the Gini-covariance (Yitzhaki and Schechtman [45]). ∎
@@
The two first terms of the previous relation correspond to the two Gini-covariances (up to ): with the rank of and conversely with the ranks of . The two last terms are (up to ) generalized Gini Mean Difference, which correspond to the intrinsic variability of and respectively. Everything happens as if the generalized Gini prametric measured on points belonging to different groups captures the two Gini-covariances between groups in excess of the covariances within groups. In this respect, it could be included in the family of distances (F3) based on scalar products.
@@ 4.1 Gini KNN
The proof is exactly the same as the one of Cover [9]. The convergence is true for any type of metric used and the bound on the error rate does not change. ∎
@@ 4.2 Gini K-means
The Gini K-means algorithm based on the generalized Gini prametric may be implemented if two conditions are met. First, convergence must be ensured for any given dataset and hyper-parameters .
Proposition 4.
The Gini K-means algorithm based on the generalized Gini prametric is convergent whenever rank vectors stay constant, for all .
@@
See the Supplementary Materials [27]. ∎
Second, since convergence is always guaranteed, an optimal value may be derived. Setting training and testing data as , a grid search may be conducted as follows. The Gini K-means algorithm is trained on and the silhouette criterion [31] is computed for each value of (not exceeding 6, otherwise too much weight is placed on minimal values of ). The optimal hyper-parameter is the value that maximizes the silhouette score. Once is determined, the labels of the testing data can be inferred.
@@
Finally, other clustering techniques relying on the Euclidean distance, close to K-means, may be adapted to the generalized Gini prametric, such as the agglomerative clustering (see subsection 5.3). An illustration is depicted in Figure 2 on the iris dataset.444 See Github.
@@ 5 Experiments
K-means and KNN classifications have been performed adapting the scikit-learn library to the generalized Gini prametric. An additional illustration is also conducted with agglomerative clustering. These algorithms are compared on the basis of 12 distances issued from the 8 families (described in Table 1): L Minkowski distance (Euclidean, Minkowski, Manhattan), L1 distance (Canberra), Inner product distance (Cosine), Squared Chord distance (Hellinger distance), Squared L2 distance (Pearson ), Shannon entropy distance (Jensen-Shannon), Vicissitude distance (Vicis Symmetric), and other distances (Hassanat). The comparison is performed over 16 datasets of UCI repository555UCI datasets repository link., without normalization in order to observe the impact of outliers (if present) on the performance. Following the study of Alfeilat et al. [1], the 16 datasets have been chosen with regard to the excellent performance of some distances robust to outliers, in particular the Hassanat distance. The datasets are described in Table 2. The detailed metrics of the Gini K-means and Gini KNN experiments, analyzed in the following subsections, are relegated in Supplementary Materials [27].
@@ 5.1 KNN Experiments
The generalized Gini prametric is implemented with the supervised KNN algorithm using the scikit-learn library. To compare the 12 distances listed above, precision and recall metrics are used on the same UCI datasets described in Table 2. To identify the optimal number of neighbors, is looped from to for each distance. A grid search is then performed to select the optimal hyper-parameter of the generalized Gini prametric that maximizes the precision. Then, recall and precision are computed and a 3-fold cross-validation is performed to validate the results (see Tables 1-10 in Supplementary Materials for experiments with and without noise [27]).
@@
Wilcoxon test. To better highlight the performance differences between the KNN methods, Table 5 shows the Wilcoxon signed-rank test results, assessing whether there are statistically significant differences in recall and precision distributions across the 16 datasets. No statistical difference is observed between the generalized Gini parametric (or Gini) and Hassanat distance. However, a significant statistical difference is found between the generalized Gini parametric and cosine (and at the 10% significance level for Gini vs. cosine).
| Comparison | Precision | Recall |
| Gini prametric vs. Hassanat | 0.1944 | 0.4842 |
| Gini prametric vs. Cosine | 0.0934 | 0.0822 |
| Generalized Gini prametric vs. Hassanat | 0.9441 | 0.5393 |
| Generalized Gini prametric vs. Cosine | 0.0113 | 0.0020 |
With 5% and 10% Noise. The experiment involves 5% and 10% noise into the data. Specifically, Gaussian noise with a mean of 0 and a standard deviation of 1 is added. This setup is designed to evaluate the robustness of the KNN algorithms under less favorable conditions, where the generalized Gini prametric is expected to outperform the other methods. In the first case, the Gini KNN achieves the highest average precision (11 top-ranked positions out of 16) and recall (10 top-ranked positions). In the 10% contamination case, the performances are much more variable. The generalized Gini prametric achieves only 6 best positions for precision and 5 for recall. Then, it comes Hassanat, Minkowski and Cosine (2 top positions for precision), and Hassanat again for recall (3 top positions, see Tables 3-10 in Supplementary Materials [27]).
Time complexity. The time complexity of the generalized Gini prametric is for a matrix of size . To compare the running time of Gini KNN and standard KNN (Minkowski with ), the Fashion MNIST dataset is used [42]. The training set consists of 49,000 images and the test set consists of 21,000 images, each of size . The Gini KNN () achieves a macro F-measure nearly identical to the Minkowski KNN (0.86 vs. 0.85 respectively). The running time is approximately 3.5 minutes on an RTX 8000 GPU for both methods. Both distances are precomputed (using the KNN implementation from the sklearn library), and only a single loop is used in the computation of each distance for a fair comparison.